I do not know how to solve system equations using substitions with fractions. Do i multiply by a common denominator?
here is an example if you could solve it step by step for me:
7x-y=10
x=1/3y-2
i know the answer (4,18) but not how to get there. if you can help it would be greatly appreciated thank you.

Answer by ptaylor(2198)

I'm confused. In the second equation is it 1/(3y) or (1/3)y. I will assume that it's the latter, so:
7x-y=10-------------------------eq1
x=(1/3)y-2---------------------eq2
If we get rid of the fraction in eq2, then you may not have to worry about it, so:
Multiply each term in eq2 by 3 and we get
3x=y-6 add 6 to each side
3x+6=y----eq2a
now substitute 3x+6 for y in eq 1
7x-(3x+6)=10
7x-3x-6=10 add 6 to each side
7x-3x-6+6=10+6 collect like terms
4x=16
x=4
substitute x=4 into eq2a and we get:
3*4+6=y
y=18
Hope this helps---ptaylor